Suppose that Cov(X,Y ) = 0.9 and Cov(X,Z) = −0.7.

(a) What is Cov(X,Y + Z)?

b) What is Cov(3X,−2Y )?
Answer #1

a) cov(x, y + z)=cov(x,y)+cov(x,z)=0.9 -0.7=0.2 b)cov(3x,–2y) =(3)(-2)cov(x,y),sin ce cov(ax,by)= ab cov(x, y) =-6(0.9) =-5.4